Calculate the molality of 9.46 g ethylene glycol (C2H6O2) in 360 g water. 




1
Expert's answer
2021-02-11T04:49:09-0500

b(ethylene glycol) = n(ethylene glycol) / m(H2O, kg)

n(ethylene glycol) = m(ethylene glycol) / M(ethylene glycol) = 9.46 / 62 = 0.153 (mol)

b(ethylene glycol) = 0.153 / 0.36 = 0.425 (mol/kg)

Answer: b = 0.425 (mol/kg).